A boy throws a ball vertically upwards and catches it back in 10s.calculate the velocity with which the ball is thrown and the height attained by it if value of g is 10ms^2
1
Expert's answer
2017-08-25T09:28:06-0400
The time of flight is
T=2t=10 s.
The height is
h=(gt^2)/2=〖10(10/2)〗^2/2=125 m.
The initial velocity is
v=gt=10(10/2)=50 m/s.
Answer: 50 m/s; 125 m.